OpenStudy (anonymous):

Simplify one over two(g − 16)

OpenStudy (anonymous):

you mean like 

OpenStudy (anonymous):

\[1\over2(g-16)\]just distribute the 2 in the denominator

OpenStudy (anonymous):

and in my opinion, that's already as simplified as it gets lol

OpenStudy (anonymous):

\[\frac{ 1 }{2 }(g-16)\] \[\frac{ 1 }{ 2 }g-8\] (16/2=8)